# A Survey on Medical Large Language Models: Technology, Application, Trustworthiness, and Future Directions

### v1

@@ IV-A1 General Algorithm
The term “Chain-of-Thought” refers to a cognitive process where an individual mentally links a series of thoughts, ideas, or reasoning steps to solve a problem, make a decision, or understand a concept. This can include breaking down a question into sub-problems, considering possible solutions, eliminating incorrect options based on available knowledge, and justifying the final answer. By generating intermediate steps and explanations, LLM models become more transparent in their decision-making, facilitating understanding and trust.
@@ IV-A2 Specific Reasoning Techniques
In-Context Padding (ICP) [41]. ICP consists of four major steps to improve the reasoning capacity of LLM in the context of clinical environments: (1) extracting medical entities from the clinical context and the reasoning objective; (2) inferring relevant medical entities from KG; (3) concatenating the acquired knowledge seeds with the prompt; (4) generating the reasoning results as well the clinical explanation.
JMLR [42]. Unlike previous RAG methods where the retrieval model is trained separately from the LLM, JMLR [42] jointly trains the LLM and information retrieval model during the fine-tuning stage. A synchronized training mechanism is formulated to retrieve clinical guidelines and leverage medical knowledge, significantly improving the clinical reasoning capabilities of LLMs and decrease the requirements for computational resources.
@@ IV-B1 General Algorithms
  • Inference-Time KG Augmentation: In this approach, an LLM is augmented with a retrieval mechanism that fetches relevant information from a KG based on the context of the input [192, 193, 194]. The model can then use this external knowledge to inform its output [195], reducing the likelihood of hallucinations and enhancing the domain-specific accuracy of its responses [196]. This is akin to giving the model a dynamic cheat sheet of verified facts to reference during its processing.
  • Training-Time KG Augmentation: This strategy goes beyond simple retrieval by creating a more integrated system where the LLM and KG interact more intimately [197, 198]. It might involve training the LLM on tasks that explicitly leverage KG data, allowing the model to learn how to utilize structured knowledge more effectively within its natural language processing capabilities. This could include updating the model’s parameters based on KG interactions, essentially teaching it how to reason with structured data.
@@ IV-B2 Specific KG-Augmented Med-LLMs
DR.KNOWS [43]. Guided by clinical diagnostic standards, DR.KNOWS is introduced as an innovative approach to improve LLMs in generating diagnostic outcomes. This method combines the utilization of a medical knowledge graph and the Unified Medical Language System (UMLS) of the National Library of Medicine. DR.KNOWS serves as an interpretive and summarizing aide, harnessing the medical KG to untangle complex medical notions. Tt establishes a clear, justifiable route to diagnoses, facilitating the employment of AI-augmented decision support mechanisms that offer transparency in their analytical processes. Consequently, this integration advances the precision of AI diagnoses while ensuring they remain comprehensible and accountable.
KG-Rank [18]. To tackle the issues of factual inconsistencies and inherent biases in medical inquiries, the KG-Rank framework [18] is introduced as an enhancement to LLMs. This framework employs a medical knowledge graph augmented with ranking and re-ranking strategies, specifically designed to enhance the accuracy of free-text medical question-answering. Concretely, for a specific question, the triplets are initially retrieved from the medical KG to gather factual information. Subsequently, ranking algorithms are applied to refine this initial retrieval results, ordering the results in a manner that facilitates the generation of more precise responses. According to the report [18], KG-Rank pioneers the integration of ranking models with KGs in the context of medical question answering for producing elaborate answers.
MedKgConv [44]. Existing generative dialog approaches may suffer from monotonous and uninteresting conversations. To address this issue, MedKgConv [44] proposed to combine various pre-trained language models with UMLS to generate human-like conversations based on MedDialog-EN dataset. UMLS contains diverse medical-related information, i.e.. disease, symptoms, and laboratory tests. To apply semantic information from the graphs, the reasoning step is conducted over the retrieved KG by reading the triples in each graph using MedFact attention. Then, a policy network is used to effectively inject relevant entities into the response text.
ChiMed [45]. The Chinese medicine dataset is built to train the Qilin-med model, including medical question answering, plain texts, knowledge graphs, and dialogues. The knowledge graph subset contains the data from CPubMed-KG [199], 39Health-KG [200], and Xywy-KG [201]. Diverse medical information is included to ensure the KG comprehensiveness, e.g., causation, symptoms, and recommended drugs.
DISC-MedLLM [46]. To construct high-quality SFT datasets, DISC-MedLLM [46] proposed a medical KG-driven sample construction to generate accurate and truthful medical responses. A department-oriented strategy is utilized to select triples from a medical KG according to the patient queries. For each triple, GPT-3.5 is exploited to generate QA pairs in a few-shot manner.
@@ IV-C LLM-based Medical Agents
Inspired by the cutting-edge evolution of AI agents, LLMs are not just utilized for generating text but act as central systems for autonomous agent systems [202, 203, 204]. This paradigm shift represents a significant leap forward in AI capabilities, moving from passive question-answering systems to proactive, versatile agents capable of executing complex tasks.
@@ IV-C2 Specific Medical Agent
CT-Agent [47]. Motivated by the advanced LLMs and multi-agent systems, CT-Agent (Clinical Agent System) is introduced as an integrated approach to improve accessibility and utility for clinical tasks. CT-Agent leverages GPT-4 [93], multi-agent architectures, LEAST-TO-MOST [211], and ReAct [210] reasoning technology, which can boost the performance in clinical contexts for managing the entire clinical process.
AutoGen [48]. AutoGen is an open-source and customizable framework for building effective and efficient applications, including mathematics, coding, QA, and online decision-making. Based on AutoGen, users can design multi-agent applications in a convenient manner, where interactions among them are conducted to finish final tasks. AutoGen can work in various modes via the combinations of LLMs, human inputs, and tools. As stated in [48], both natural language and computer code can be applied to formulate conversation patterns. Therefore, AutoGen-based users can flexibly define agent behaviors.
ArgMed-Agents [49]. To satisfy the requirements of clinical tasks (e.g., complex reasoning and planning), a multi-agent framework called ArgMed-Agents is proposed to provide explainable clinical decision reasoning. ArgMed-Agents is able to mimic the procedure of clinical reasoning by generating explanations in a self-directed way. The core is to perform self-argumentation iterations based on a reasoning mechanism for modeling cognitive processes. Then an augmented process is taken as a directed graph to represent conflicting relationships. Then, a symbolic solver can identify rational and coherent arguments for the decision.
MAD [50]. Multi-agent debate (MAD) [50] is proposed as a useful strategy to enhance the truthfulness of LLMs, which explores various debating and prompting strategies to maintain the trade-offs among cost, time, and accuracy. MAD protocols are difficult to optimize due to their sensitivity to hyperparameters. Some insights are proposed to improve debating strategies, such as adjusting agent agreement levels.
@@ IV-D2 Specific RAG Algorithm
Clinfo.ai [51]. Clinfo.ai is an open-source WebApp that answers clinical questions based on dynamically retrieved scientific literature. The information retrieval and summarization tasks are applied to evaluate the retrieval-augmented LLM systems.
@@
Almanac [52]. Incorrect and harmful generations indeed limit the clinical applications of LLMs. To address this issue, Almanac is equipped with retrieval capabilities from curated medical resources for clinical guidelines and treatment recommendations. A group of clinicians and health care practitioners are utilized to evaluate Almanac, which makes a comparison between the responses from Almanac and standard LLMs (e.g., ChatGPT-4).
@@
BiomedRAG [53]. Different from previous retrieval-augmented LLMs, BiomedRAG [53] exploits a straightforward RAG method by directly inputting the retrieved documents into the LLM, rather than utilizing cross-attention mechanisms to encode retrieved text. BiomedRAG can effectively decrease the negative effects of noise information in retrieved documents, particularly for noise-intensive tasks. The study demonstrated the effectiveness of LLMs working as the supervised signal for the retrieval model in the biomedical domain.
@@
Self-BioRAG [54]. To improve RAG generalization for domain-specific problems, the Self-BioRAG framework [54] is introduced for biomedical text, which specializes in explanation generation, domain-specific document retrieval, and self-reflecting generated responses. 84k biomedical instruction sets are exploited to train Self-BioRAG. The study indicated that domain-specific components, such as a retriever, domain-related document corpus, and instruction sets are necessary for adhering to domain-related instructions.
@@
ECG-RAG [55]. To investigate LLMs for the ECG diagnosis, a zero-shot retrieval-augmented diagnosis technique [55] is designed to utilize the inherent encoded knowledge while infusing expert knowledge for carefully crafting prompts. Datasets of specific domain knowledge are built filled with cardiac symptoms and sleep apnea diagnosis.
@@
ChatENT [56]. Existing LLMs suffer from inherent limitations including inconsistent accuracy, specific prompting requirements, and the risk of generating harmful hallucinations [13]. A domain-specific fine-tuned model would effectively address these limitations. OHNS-relevant data is gathered from open-access internet sources. Retrieval-Augmented Language Modeling (RALM) [56] is utilized to recall this information for pre-training, which is integrated into ChatGPT to create an OHNS-specific knowledge QA platform, named as ChatENT.
@@
MIRAGE [57]. Medical Information Retrieval-Augmented Generation Evaluation (MIRAGE) [57] is constructed to evaluate the medical RAG systems, including 7,663 questions from five medical QA datasets. MEDRAG can improve the performance of various LLMs, outperforming Chain-of-Thought prompting, GPT-3.5, Mixtral, and GPT-4 .
@@
MedicineQA [58]. Due to the lack of domain-specific knowledge, it is challenging to deploy LLM models into medical scenarios. A multi-round dialogue benchmark called MedicineQA [58] is proposed to simulate the real-world medication scenarios, requiring LLMs to answer with retrieved evidence from the medicine datasets. MedicineQA contains 300 multi-round question-answering pairs. Then a Distill-Retrieve-Read framework is designed to replace previous Retrieve-then-Read, which utilizes a tool calling mechanism to formulate search queries.
@@ IV-E2 Specific Alignment Algorithm
Safety Alignment [59]. The research carries out the first safety evaluation for Med-LLMs. To this end, the study [59] summarizes the definitions of medical safety and alignment for medical AI systems. Then an evaluation dataset is constructed with harmful medical questions as indicators, which evaluates both general and medical safety and alignment of Med-LLMs.
SELF-ALIGN [60]. Due to the expensive costs of collecting human supervision and the quality issues (e.g., reliability, diversity, self-consistency, and undesirable biases), a SELF-ALIGN approach is proposed to combine principle-driven reasoning and the generative power of LLMs [60], which aims to achieve the self-alignment with minimal human supervision. SELF-ALIGN contains four steps: (1) synthetic prompt generation to improve prompt diversity; (2) human-written principles for response generation; (3) high-quality self-aligned fine-tuning; (4) a refinement step to avoid simple or indirect responses.
EGR [61]. It is effective for few-shot prompting to combine diverse techniques, which can significantly enhance the performance of LLMs. Motivated by this, the EGR (expand-guess-refine) alignment strategy [61] is proposed for medical QA as a parameter and data-efficient solution.
@@ IV-F1 General Algorithms
LLaVA. The Large Language and Vision Assistant (LLaVA) model [217] is introduced to bridge vision and language understanding through a multi-modal approach. It combines a vision encoder (CLIP [218]) with a language decoder (Vicuna [219]) to form a general-purpose MM-LLM for vision-to-language tasks. The training process involves fine-tuning the decoder on 158,000 language-vision instruction examples from the MS-COCO dataset [220].
mPLUG-OWL. mPLUG-OWL [221] is first introduced by the researchers from the Alibaba DAMO academy, which could lead to inadequate alignment due to limited parameter flexibility. mPLUG-OWL utilizes advanced pre-trained LLM LLaMA-7B [105] as the language decoder, while ViT-L/14 serves as the visual foundation model for the visual encoder to extract visual features from the input images. The ViT is initialized from the pre-trained CLIP ViT-L/14 model [218] for faster convergence.

### v2

@@ IV-A1 General Algorithm
The term ”Chain-of-Thought” refers to a cognitive process where an individual mentally links a series of thoughts, ideas, or reasoning steps to solve a problem, make a decision, or understand a concept. This can include breaking down a question into sub-problems, considering possible solutions, eliminating incorrect options based on available knowledge, and justifying the final answer. By generating intermediate steps and explanations, LLMs become more transparent in their decision-making, facilitating understanding and trust.
@@ IV-A2 Specific Reasoning Techniques
$\bullet$ In-Context Padding (ICP) [41]. ICP consists of four major steps to improve the reasoning capacity of LLM in the context of clinical environments: (1) extracting medical entities from the clinical context and the reasoning objective; (2) inferring relevant medical entities from KG; (3) concatenating the acquired knowledge seeds with the prompt; (4) generating the reasoning results as well the clinical explanation.
$\bullet$ JMLR [42]. Unlike previous RAG methods where the retrieval model is trained separately from the LLM, JMLR [42] jointly trains the LLM and information retrieval model during the fine-tuning stage. A synchronized training mechanism is formulated to retrieve clinical guidelines and leverage medical knowledge, significantly improving the clinical reasoning capabilities of LLMs and decrease the requirements for computational resources.
@@ IV-B1 General Algorithms
$\bullet$ Inference-Time KG Augmentation. In this approach, an LLM is augmented with a retrieval mechanism that fetches relevant information from a KG based on the context of the input [184, 185, 186]. The model can then use this external knowledge to inform its output [187], reducing the likelihood of hallucinations and enhancing the domain-specific accuracy of its responses [188]. This is akin to giving the model a dynamic cheat sheet of verified facts to reference during its processing.
$\bullet$ Training-Time KG Augmentation. This strategy goes beyond simple retrieval by creating a more integrated system where the LLM and KG interact more intimately [189, 190]. It might involve training the LLM on tasks that explicitly leverage KG data, allowing the model to learn how to utilize structured knowledge more effectively within its natural language processing capabilities. This could include updating the model’s parameters based on KG interactions, essentially teaching it how to reason with structured data.
@@ IV-B2 Specific KG-Augmented Med-LLMs
$\bullet$ DR.KNOWS [43]. Guided by clinical diagnostic standards, DR.KNOWS is introduced as an innovative approach to improve LLMs in generating diagnostic outcomes. This method combines the utilization of a medical knowledge graph and the Unified Medical Language System (UMLS) of the National Library of Medicine. DR.KNOWS serves as an interpretive and summarizing aide, harnessing the medical KG to untangle complex medical notions. Tt establishes a clear, justifiable route to diagnoses, facilitating the employment of AI-augmented decision support mechanisms that offer transparency in their analytical processes. Consequently, this integration advances the precision of AI diagnoses while ensuring they remain comprehensible and accountable.
$\bullet$ KG-Rank [18]. To tackle the issues of factual inconsistencies and inherent biases in medical inquiries, the KG-Rank framework [18] is introduced as an enhancement to LLMs. This framework employs a medical knowledge graph augmented with ranking and re-ranking strategies, specifically designed to enhance the accuracy of free-text medical question-answering. Concretely, for a specific question, the triplets are initially retrieved from the medical KG to gather factual information. Subsequently, ranking algorithms are applied to refine this initial retrieval results, ordering the results in a manner that facilitates the generation of more precise responses. According to the report [18], KG-Rank pioneers the integration of ranking models with KGs in the context of medical question answering for producing elaborate answers.
$\bullet$ MedKgConv [44]. Existing generative dialog approaches may suffer from monotonous and uninteresting conversations. To address this issue, MedKgConv [44] proposed to combine various pre-trained language models with UMLS to generate human-like conversations based on MedDialog-EN dataset. UMLS contains diverse medical-related information, i.e.. disease, symptoms, and laboratory tests. To apply semantic information from the graphs, the reasoning step is conducted over the retrieved KG by reading the triples in each graph using MedFact attention. Then, a policy network is used to effectively inject relevant entities into the response text.
$\bullet$ ChiMed [45]. The Chinese medicine dataset is built to train the Qilin-med model, including medical question answering, plain texts, knowledge graphs, and dialogues. The knowledge graph subset contains the data from CPubMed-KG [191], 39Health-KG [192], and Xywy-KG [193]. Diverse medical information is included to ensure the KG comprehensiveness, e.g., causation, symptoms, and recommended drugs.
$\bullet$ DISC-MedLLM [46]. To construct high-quality SFT datasets, DISC-MedLLM [46] proposed a medical KG-driven sample construction to generate accurate and truthful medical responses. A department-oriented strategy is utilized to select triples from a medical KG according to the patient queries. For each triple, GPT-3.5 is exploited to generate QA pairs in a few-shot manner.
@@ IV-C LLM-based Medical Agents
Inspired by the cutting-edge evolution of AI agents, LLMs are not just utilized for generating text but act as central systems for autonomous agent systems [194, 195, 196]. This paradigm represents a significant leap forward in AI capabilities, moving from passive question-answering systems to proactive, versatile agents capable of executing complex tasks.
@@ IV-C2 Specific Medical Agent
Here, we introduce some LLM-based medical agents. These agents should complement rather than replace human healthcare providers, ensuring that the nuanced understanding and empathy that humans bring to patient care are not lost.
$\bullet$ CT-Agent [47]. Motivated by the advanced LLMs and multi-agent systems, CT-Agent (Clinical Agent System) is introduced as an integrated approach to improve accessibility and utility for clinical tasks. CT-Agent leverages GPT-4 [203], multi-agent architectures, LEAST-TO-MOST [204], and ReAct [202] reasoning technology, which can boost the performance in clinical contexts for managing the entire clinical process.
$\bullet$ AutoGen [48]. AutoGen is an open-source and customizable framework for building effective and efficient applications, including mathematics, coding, QA, and online decision-making. Based on AutoGen, users can design multi-agent applications in a convenient manner, where interactions among them are conducted to finish final tasks. AutoGen can work in various modes via the combinations of LLMs, human inputs, and tools. As stated in [48], both natural language and computer code can be applied to formulate conversation patterns. Therefore, AutoGen-based users can flexibly define agent behaviors.
$\bullet$ ArgMed-Agents [49]. To satisfy the requirements of clinical tasks (e.g., complex reasoning and planning), a multi-agent framework called ArgMed-Agents is proposed to provide explainable clinical decision reasoning. ArgMed-Agents is able to mimic the procedure of clinical reasoning by generating explanations in a self-directed way. The core is to perform self-argumentation iterations based on a reasoning mechanism for modeling cognitive processes. Then an augmented process is taken as a directed graph to represent conflicting relationships. Then, a symbolic solver can identify rational and coherent arguments for the decision.
$\bullet$ MAD [50]. Multi-agent debate (MAD) [50] is proposed as a useful strategy to enhance the truthfulness of LLMs, which explores various debating and prompting strategies to maintain the trade-offs among cost, time, and accuracy. MAD protocols are difficult to optimize due to their sensitivity to hyperparameters. Some insights are proposed to improve debating strategies, such as adjusting agent agreement levels.
@@ IV-D2 Specific RAG Algorithm
We summarize some RAG applications in the medical domain. By integrating external knowledge with generative models, RAG can provide more accurate, up-to-date, and contextually relevant information compared to traditional methods.
@@
$\bullet$ Clinfo.ai [51]. Clinfo.ai is an open-source WebApp that answers clinical questions based on dynamically retrieved scientific literature. The information retrieval and summarization tasks are applied to evaluate the retrieval-augmented LLM systems.
@@
$\bullet$ Almanac [52]. Incorrect and harmful generations indeed limit the clinical applications of LLMs. To address this issue, Almanac is equipped with retrieval capabilities from curated medical resources for clinical guidelines and treatment recommendations. A group of clinicians and health care practitioners are utilized to evaluate Almanac, which makes a comparison between the responses from Almanac and standard LLMs (e.g., ChatGPT-4).
@@
$\bullet$ BiomedRAG [53]. Different from previous retrieval-augmented LLMs, BiomedRAG [53] exploits a straightforward RAG method by directly inputting the retrieved documents into the LLM, rather than utilizing cross-attention mechanisms to encode retrieved text. BiomedRAG can effectively decrease the negative effects of noise information in retrieved documents, particularly for noise-intensive tasks. The study demonstrated the effectiveness of LLMs working as the supervised signal for the retrieval model in the biomedical domain.
@@
$\bullet$ Self-BioRAG [54]. To improve RAG generalization for domain-specific problems, the Self-BioRAG framework [54] is introduced for biomedical text, which specializes in explanation generation, domain-specific document retrieval, and self-reflecting generated responses. 84k biomedical instruction sets are exploited to train Self-BioRAG. The study indicated that domain-specific components, such as a retriever, domain-related document corpus, and instruction sets are necessary for adhering to domain-related instructions.
@@
$\bullet$ ECG-RAG [55]. To investigate LLMs for the ECG diagnosis, a zero-shot retrieval-augmented diagnosis technique [55] is designed to utilize the inherent encoded knowledge while infusing expert knowledge for carefully crafting prompts. Datasets of specific domain knowledge are built filled with cardiac symptoms and sleep apnea diagnosis.
@@
$\bullet$ ChatENT [56]. Existing LLMs suffer from inherent limitations including inconsistent accuracy, specific prompting requirements, and the risk of generating harmful hallucinations [13]. A domain-specific fine-tuned model would effectively address these limitations. OHNS-relevant data is gathered from open-access internet sources. Retrieval-Augmented Language Modeling (RALM) [56] is utilized to recall this information for pre-training, which is integrated into ChatGPT to create an OHNS-specific knowledge QA platform, named as ChatENT.
@@
$\bullet$ MIRAGE [57]. Medical Information Retrieval-Augmented Generation Evaluation (MIRAGE) [57] is constructed to evaluate the medical RAG systems, including 7,663 questions from five medical QA datasets. MEDRAG can improve the performance of various LLMs, outperforming Chain-of-Thought prompting, GPT-3.5, Mixtral, and GPT-4 .
@@
$\bullet$ MedicineQA [58]. Due to the lack of domain-specific knowledge, it is challenging to deploy LLM models into medical scenarios. A multi-round dialogue benchmark called MedicineQA [58] is proposed to simulate the real-world medication scenarios, requiring LLMs to answer with retrieved evidence from the medicine datasets. MedicineQA contains 300 multi-round question-answering pairs. Then a Distill-Retrieve-Read framework is designed to replace previous Retrieve-then-Read, which utilizes a tool calling mechanism to formulate search queries.
@@ IV-E2 Specific Alignment Algorithm
Achieving human alignment in medical LLMs is crucial because it directly impacts patient safety, trust in technology, and the overall effectiveness of healthcare delivery. Here are some human alignment algorithms, which can be approached in the development and deployment of medical LLMs.
$\bullet$ Safety Alignment [59]. The research carries out the first safety evaluation for Med-LLMs. To this end, the study [59] summarizes the definitions of medical safety and alignment for medical AI systems. Then an evaluation dataset is constructed with harmful medical questions as indicators, which evaluates both general and medical safety and alignment of Med-LLMs.
$\bullet$ SELF-ALIGN [60]. Due to the expensive costs of collecting human supervision and the quality issues (e.g., reliability, diversity, self-consistency, and undesirable biases), a SELF-ALIGN approach is proposed to combine principle-driven reasoning and the generative power of LLMs [60], which aims to achieve the self-alignment with minimal human supervision. SELF-ALIGN contains four steps: (1) synthetic prompt generation to improve prompt diversity; (2) human-written principles for response generation; (3) high-quality self-aligned fine-tuning; (4) a refinement step to avoid simple or indirect responses.
$\bullet$ EGR [61]. It is effective for few-shot prompting to combine diverse techniques, which can significantly enhance the performance of LLMs. Motivated by this, the EGR (expand-guess-refine) alignment strategy [61] is proposed for medical QA as a parameter and data-efficient solution.
@@ IV-F1 General Algorithms
$\bullet$ PaLM-E [210]. PaLM-E is part of Google’s Pathways initiative, which aims to create AI systems that can perform a wide range of tasks across different modalities. It builds upon the success of PaLM, a large-scale text-only language model, by integrating multimodal capabilities. PaLM-E incorporates both textual and visual embeddings, allowing it to understand and generate content in response to images and text prompts. It can perform image captioning, visual question answering, and even complex reasoning about visual scenes, all without requiring extensive fine-tuning on task-specific datasets.
$\bullet$ LLaVA [211]. The Large Language and Vision Assistant (LLaVA) model is introduced to bridge vision and language understanding through a multi-modal approach. It combines a vision encoder (CLIP [212]) with a language decoder (Vicuna [213]) to form a general-purpose MM-LLM for vision-to-language tasks. The training process involves fine-tuning the decoder on 158,000 language-vision instruction examples from the MS-COCO dataset [214].
$\bullet$ mPLUG-OWL [215]. mPLUG-OWL is first introduced by the researchers from the Alibaba DAMO academy, which could lead to inadequate alignment due to limited parameter flexibility. mPLUG-OWL utilizes advanced pre-trained LLM LLaMA-7B [115] as the language decoder, while ViT-L/14 serves as the visual foundation model for the visual encoder to extract visual features from the input images. The ViT is initialized from the pre-trained CLIP ViT-L/14 model [212] for faster convergence.